OpenStudy (anonymous):

The mass of the wedge (movable wedge) is 7.89 kg the coefficient of friction is U b/w the 0.5 kg and the wedge for which the wedge just begins to move to the right when released from rest

OpenStudy (anonymous):



OpenStudy (anonymous):

@demitris

OpenStudy (anonymous):

What's the question here?

OpenStudy (anonymous):

we have to find U?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Friction#Dry_friction Force down=\[0.4g\]=tension upwards Normal force=\[0.5g \cdot \cos(37)\] Tension upwards=normal force * U \[0.5g \cdot \cos(37) U=0.4g\]

OpenStudy (anonymous):

I'm oversimplifying, aren't I?

OpenStudy (anonymous):

4 - T = 0.4a T - mg sin 37 - U mg cos 37 = ma

OpenStudy (anonymous):

Shuld it be like this ?

OpenStudy (anonymous):

The only net force on the system is down, and since I"m assuming the block is resting on a (frictionless) table or something so it can't move in the y direction, that means that the forces on the system in the horizontal direction must cancel. If there's no net force, that means the momentum is constant -- it follows that the wedge will begin to move as soon as the masses do. Therefore, you just need to find the coefficient of friction such that the masses just begin to slide.

OpenStudy (anonymous):

u mean .... mg sin 37 = U mg cos 37

OpenStudy (anonymous):

Nope. Just equate the forces tugging at each end of the string (including friction)

OpenStudy (ghazi):

 0.5g sin 37= T- U*0.5*g cos 37...........(i) T= 0.4*g